OpenStudy (anonymous):

Sarah used her calculator to find sin125. She wrote down sin125 = -.57. How could Sarah recognize that her answer is incorrect? Please someone help me with this problem. I have no idea how to do it and I've been staring at it for the last 2 hours.

OpenStudy (anonymous):

@Science_ALLY

OpenStudy (anonymous):

@Science_ALLY any help is appreciated at this point

OpenStudy (science_ally):

im sowwy idk anything about sin ....I cant help but I might b able 2 find sum1

OpenStudy (anonymous):

@Science_ALLY if you could find anyone that would be great

OpenStudy (science_ally):

kk , i'll try ... give me a few sec

OpenStudy (science_ally):

found someone @danish071996 , he's going 2 try

OpenStudy (anonymous):

@Science_ALLY Great! Thanks so much

OpenStudy (science_ally):

well I didn't really help much ... but u welcome

rvc (rvc):

hey do u know what is \[\sin(\Pi-x)=?\]

rvc (rvc):

@ivanmorozov pls ans

OpenStudy (anonymous):

@rvc no I haven't seen that before

rvc (rvc):

oh sin (180-x)=sin x

OpenStudy (anonymous):

@rvc what does (180-x)=sin x mean?

rvc (rvc):

ohh put x = 125

OpenStudy (anonymous):

@rvc is the answer 0.82?

rvc (rvc):

180-125=55 sin55=?

OpenStudy (anonymous):

@rvc sin55= 0.819

rvc (rvc):

sin55=0.819 = 0.82

OpenStudy (anonymous):

So how could Sarah figure out that her answer is incorrect? @rvc

rvc (rvc):

its wrong sin55=0.82 x=55\[\sin(55)=\sin(x)\rightarrow x=55\]

OpenStudy (anonymous):

Oh I get it now. Thank you so much! @rvc

rvc (rvc):

no thanks please @ivanmorozov all the best

OpenStudy (anonymous):

Where is the toilet?